OpenStudy (radvedder):

Solving inequalities??? help please!! 1/4m is less than or equal to -17

OpenStudy (radvedder):



OpenStudy (anonymous):

Solving for m right? Just solve it like a regular equation, pretend the inequality sign is just an "equal" sign

OpenStudy (radvedder):

but i have to multiply or divide from each side and i don't know if i multiply by -17 or divide by 1/4

OpenStudy (johnweldon1993):

Hint* \[\large \frac{1}{4}m \le -17\] Since we want to solve for 'm' we need to get rid of that 1/4 being multiplied to it somehow

OpenStudy (radvedder):

so multiply that on each side?

OpenStudy (johnweldon1993):

Multiply what?

OpenStudy (radvedder):

1/4

OpenStudy (johnweldon1993):

Well lets see what happens when we do that...if we multiply both sides by 1/4 we have \[\large \frac{1}{4} \times \frac{1}{4}m \le -17 \times \frac{1}{4}\] \[\large \frac{1}{16}m \le \frac{-17}{4}\] That didnt help anything...where did we go wring?

OpenStudy (anonymous):

You want to cancel out the factor....how do you do that? *hint do not multiply 1/4

OpenStudy (radvedder):

I don't know I suck at math. would i multiply the -17? or divide?

OpenStudy (anonymous):

How do you get 1/4 to = 1?

OpenStudy (anonymous):

what do you need to multiply to 1/4?

OpenStudy (radvedder):

multiply -17 to 1/4

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }\times \frac{ ? }{ 1 }=1\]

OpenStudy (johnweldon1993):

Here's a big hint \[\large \frac{1}{b}x \times b = x\]

OpenStudy (anonymous):

Nope, multiplying -17 to 1/4 just gives you -17/4 and that does not equal to 1

OpenStudy (radvedder):

okay, we didn't do it like that in class though. i might not have explained it right, i don't know lol. i'll just go in and ask for help in the morning. but thank you guys for trying to help me

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }\times \frac{ 4 }{ 1 }=1\]

OpenStudy (anonymous):

Do you see how that would work?

OpenStudy (johnweldon1993):

Well here, [\large \frac{1}{4}x \le -17\] If we multiply both sides by 4...we have \[\large \cancel{4} \times \cancel{\frac{1}{4}}x \le -17 \times 4\] Which gives \[\large x \le -17 \times 4\]

OpenStudy (anonymous):

You're solving for m so that means you need to get m alone

OpenStudy (radvedder):

for the answer it's suppose to be something like x < 15 for example or something like that. like a variable and a less than or equal to sign and a number

OpenStudy (johnweldon1993):

Sorry for the first line not coming out correctly up there

OpenStudy (anonymous):

To get m alone, that means you need to move that fraction to the other side

OpenStudy (anonymous):

When multiplying fractions, you always multiply the inverse of that fraction (meaning the number on top and the number on the bottom switch)

OpenStudy (radvedder):

okay thank you!

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }m \le-17\] \[\left( \frac{ 4 }{ 1 } \right)\frac{ 1 }{ 4 }m \le-17\left( \frac{ 4 }{ 1 } \right)\] (multiply 4/1 to each side) then \[m \le-68\]